OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=-2x+7 y=1/3x ( ? , ? )

OpenStudy (anonymous):

@texasman21

OpenStudy (anonymous):

medal and fan

OpenStudy (anonymous):

@mathmate @TwiztTiez

OpenStudy (twizttiez):

What's up?

OpenStudy (anonymous):

i need help.

OpenStudy (anonymous):

@Demonx341

OpenStudy (anonymous):

@iGreen

OpenStudy (mathmate):

Is the second equation y=(1/3)x or y=1/(3x) ? In either case, substitute y into the first equation and solve for x.

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

y=1 over 3

OpenStudy (mathmate):

y=(1/3)*x ? Whenever you have y= and y= for both equations, just equate the right hand sides and solve for x. In this case, -2x+7=(1/3)x and solve for x. After that, substitute the value of x into either equation and solve for y.

OpenStudy (anonymous):

^ is what i would do :P

OpenStudy (anonymous):

in this case hoe do i find x?

OpenStudy (anonymous):

how i meant. lol

OpenStudy (mathmate):

Recall cross multiply, but don't forget to put parentheses around the left-hand side before you multiply.

OpenStudy (anonymous):

i hate algebra.

OpenStudy (mathmate):

You have (-2x+7)=x/3 Multiply both sides by 3 and solve for x.

OpenStudy (anonymous):

i got 1. x=-6

OpenStudy (mathmate):

Multiply by 3: 3(-2x+7)=x -6x+21=x 21=7x x=3 Now can you put x=3 into one of the two original equations and find y?

OpenStudy (anonymous):

ohh. i forgot to multiply 7. ;p

OpenStudy (anonymous):

can u explain to me how to find y?